# Test-Time Degradation Adaption for Open-Set Image Restoration

### v3

@@ 3.3 Test-time Degradation Adapter
The essence of OIR lies in addressing the unidentified distribution shifts between test and training data, while TTA recently emerges as an effective methodology to address this inherent disparities. With the motivations, TDA is presented for adapting PDM to the downstream shifted distribution of degradations during the test phase.
TDA employs a four-layer convolution neural network as the adapter $\phi$, which is optimized for aligning the domain of generated images $\hat{x}^{g}_{0}$, $\hat{x}^{r}_{0}$ to that of the test degraded image $y$. Note that $\hat{x}^{g}_{0}$, $\hat{x}^{r}_{0}$ are the estimated output of $x^{g}_{t}$, $x^{r}_{t}$, respectively. To this end, we introduce adversarial training which has shown to be effective in aligning the source and target domains, and the translated samples usually capture rich visual characteristics inherent to the target domain. Therefore, TDA aligns $\phi([\hat{x}^{g}_{0},\hat{x}^{r}_{0}])$ to $y$ which involves significant degradations through
$\mathcal{L}_{adv}=\log(1-\mathcal{D}(\phi([\hat{x}^{g}_{0},\hat{x}^{r}_{0}])),$ (5)
@@
$\mathcal{L}_{dis}=-\log\mathcal{D}([y,y])-\log(1-\mathcal{D}(\phi([\hat{x}^{g}_{0},\hat{x}^{r}_{0}])).$ (6)
Meanwhile, TDA introduces a reconstruction loss to maintain the content consistency before and after alignment, i.e.,
$\mathcal{L}_{rec}=\parallel[\hat{x}^{g}_{0},\hat{x}^{r}_{0}]-\phi([\hat{x}^{g}_{0},\hat{x}^{r}_{0}])\parallel_{2}^{2}.$ (7)
Besides, the perceptual loss is adopted to remain the semantic consistency before and after alignment, i.e.,
$\mathcal{L}_{pec}=\parallel\mathcal{V}([\hat{x}^{g}_{0},\hat{x}^{r}_{0}])-\mathcal{V}(\phi([\hat{x}^{g}_{0},\hat{x}^{r}_{0}]))\parallel_{2}^{2},$ (8)
@@
where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are the optimization weights. Through iterative optimizations during the denoising process, $\phi$ progressively translate $\hat{x}^{g}_{0}$, $\hat{x}^{r}_{0}$ to the domain of $y$. Note that $\hat{x}^{r}_{0}$ is introduced to prevent $\phi$ from the trivial solution of identity mapping.
@@ 3.4 Adapter-guided Image Restoration
With the adapter $\phi$, $\hat{x}^{g}_{0}$ could be translated to the degraded domain of $y$ through $\phi(\hat{x}^{g}_{0})$, which could be further supervised towards the $y$, and in turn produce $\mathcal{G}$ to update the generated images $\hat{x}^{g}_{0}$ to the desired clean output $x^{g}_{0}$. AIR is devised to perform this supervised updating process.
Existing studies [5] have show that the denoising process exhibits an intriguing temporal dynamic, i.e., $\hat{x}^{g}_{0}$ is generated from the unrecognizable contents to the perceptually rich contents, and finally the imperceptible details. In other words, the contents to be restored are varying as the timesteps during the denoising process. To take full advantages of PDM, AIR empirically divides the denoising process into three stages, and guides the restoration through different strategies at different stages.
In the first stage where $t\sim T$, the images have a lower signal-to-noise ratio, where the contents are unrecognizable and only contains some global information such as colors, layouts distributed throughout image. Therefore, AIR applied mse loss on the spatial pixels and color channels, i.e.,
@@
where $\gamma_{1}$ is the loss weight, $\hat{x}_{0}$ is the estimated output of $x_{t}\in\{x^{g}_{t},x^{r}_{t}\}$. AIR also use this guidance on the referred image $x^{r}_{t}$ so that it embraces similar global properties as the $y$, since it is beneficial to the optimization of adapter. Besides, a linear warmup is adopted on $\gamma_{1}$.
In the second stage where $t\sim T/2$, PDM prefers to generate the perceptually rich contents, and thus AIR focuses more on restoring the perceptual contents of $y$. To this end, the perceptual loss is adopted, i.e.,
@@
Meanwhile, the adversarial loss is also introduced to boost the perceptual quality, i.e.,
@@
-- $\log\mathcal{D}([\hat{x}^{g}_{0},\hat{x}^{r}_{0}]-\phi([\hat{x}^{g}_{0},\hat{x}^{r}_{0}])).$
Through this loss, $\hat{x}^{g}_{0}$ will be further updated toward the clean image whose degraded version matches $y$, by aligning the distribution of information loss. Besides, the mse loss is also employed to remain the global contents, i.e.,
@@
In the third stage where $t\sim 0$, the perceptually rich contents are already prepared in $\hat{x}^{g}_{0}$, AIR only required to recover the imperceptible details. In other words, $\hat{x}^{g}_{0}$ is very close to the natural image, and thus some extra priors could be introduced to regulate the final output. In consideration of generality, we only introduce the total variation loss, i.e., $\mathcal{G}=\gamma_{5}*TVLoss(\hat{x}^{g}_{0})$, where $\gamma_{5}$ is the loss weight. Note that although the third stage mainly affects the imperceptible details, the pixel values still fluctuate, and thus the guidance in the second stage is still used for a better numerical fidelity.
@@ 4 Experiments
4.1 Experimental Settings
To evaluate the effectiveness of TAO on addressing OIR, we conduct experiments on multiple tasks of image restoration with exactly the same settings, except for the loss weights ($\lambda_{1-3}$, $\gamma_{1-5}$) and the scale of guidance $s$, which are fine-tuned for different types of degradation to obtain the best results. In experiments, we employ an unconditional image diffusion model [8] pretrained on ImageNet [7] as PDM, and sets the timestep as $T=1000$, which is further divided into three stages in a heuristic way, i.e., the first stage 999-600, the second stage 600-50, and the third stage 50-0. Since PDM is pre-trained on large-scale nature images for high-quality image generation, any degradation in the test image could be regraded as the unforeseen. Both the adapter and discriminators are four-layer convolutional networks, and optimized once at each denoising timestep through Adam optimizer with default learning rate of 1e-3. The first nine layers of pretrained VGG-16 [35] network are used to extract the semantic feature maps for calculating the perceptual loss. For evaluation, considering the specificity of our methods, i.e., adaption on single test sample, we mainly compare it with those task-specific zero-shot methods, and some classic learning-based methods. To access their performance, PSNR and SSIM metrics are employed, and all experiments are conducted through PyTorch framework on Ubuntu20.04 with GeForce RTX 3090 GPUs.
4.2 Experimental Results
In this section, we evaluate our method on image restoration tasks of image dehazing, low-light image enhancement, and image denoising.
Table 1: Quantitative results of image dehazing on HSTS from RESIDE dataset. Our method outperforms the zero-shot methods, and obtains comparable results with the classic methods through supervised learning.
@@ 4.2 Experimental Results
PSNR $\uparrow$ 23.96 18.40 19.15 21.69 18.42 15.75 17.41 18.18 18.22 21.24 21.38
SSIM $\uparrow$ 0.902 0.826 0.860 0.868 0.854 0.783 0.833 0.814 0.832 0.835 0.859
@@
PSNR $\uparrow$ 13.39 17.87 18.03 14.85 17.15 15.33 15.71 11.24 15.75 16.28
SSIM $\uparrow$ 0.623 0.699 0.809 0.815 0.720 0.763 0.765 0.543 0.665 0.766
@@
PSNR $\uparrow$ 26.35 28.63 28.71 27.17 27.58 28.17
SSIM $\uparrow$ 0.757 0.795 0.864 0.846 0.808 0.827
Image Dehazing aims to remove the haze and boost the visibility of hazy image, which has spawned many methods in the past decades. In experiments, we introduce HSTS dataset from RESIDE [22] for evaluations. To be specific, RESIDE is a large scale haze image dataset, and HSTS is one of the test subsets, which contains 10 synthetic and 10 real-world hazy images. Before the evaluation, we first center crop the images along the shorter edge, and then resize them to match the image size of the PDM.
For comprehensive comparisons, we compare TAO with 10 representative methods, which are specially designed for image dehazing, and could be roughly divided into two groups, i.e., classic learning-based and zero-shot methods. Specifically, the classic learning-based methods are DehazeNet (DEN) [3], MSCNN (MSN) [32], AOD-Net (AOD) [21], and CAP [47]. The zero-shot methods are DCP [14], BCCR [29], GRM [4], NLD [2], DDIP [11] and YOLY [24].
The quantitative results are presented in Tab.1, from which one could see that our method achieves comparable even better results than those task-specific methods. To be specific, our method obtains the best performance in zero-shot methods, and outperforms YOLY and DCP with 0.14/0.024 and 2.96/0.005 in PSNR/SSIM metrics, respectively. Meanwhile, even compared to the classic learning-based methods, our method still outperforms some of them, and achieves comparable performance with the other. For instance, our method obtains 2.98/0.033 higher PSNR/SSIM value than MSN, 2.23 higher PSNR value than AOD. Although it is lower than CAP and DEN, the performance gaps are not huge, e.g., 0.31/0.009 lower than CAP and 2.58/0.043 lower than DEN in PSNR/SSIM metrics. In addition, we show and analyze their qualitative results in Fig.3, which show that our method achieves better fidelity and realness.
Figure 3: Qualitative results of image dehazing, from which one could see that existing methods excessively dehazing resulting in darkening and/or artifacting of the image. In contrast, our method obtains clearer results which are closer to the natural ground truth.
Figure 4: Qualitative results of low-light image enhancement, from which one could see that our results are not as smooth as the MBLLEN nor as dark as ZDEC. Although there is a slight color bias from the ground truth, it achieves a rational lighting of the dark images.
Low-light Image Enhancement aims at improving the perception of the image captured in the environment with poor illumination. To evaluate our method on this challenge task, we introduce the test subset from LOL [41] dataset. LOL includes 485 training and 15 test image pairs of low- and normal-light, which contain noises produced during the photo capture process. Since the images have a resolution of 400$\times$600, we first center crop the image along the shorter edges, and then resize them to be applicable to the PDM.
For comprehensive comparisons, we compare our method with nine low-light image enhancement methods, which are divided into two categories, i.e., classical learning-based and zero-shot methods. To be specific, the classical learning-based methods are LightenNet (LNet) [26], Retinex-Net (RNet) [40], MBLLEN [28], and EnlightenGAN (EGAN) [16]. The zero-shot methods are ExCNet [43], Zero-DCE (ZDEC) [13], Zero-DCE++ (ZDEC+) [27], RRDNet (RRDN) [46], and GDP [9]. The above methods except GDP are specially designed for low-light image enhancement, while our method is devised for more general purpose of OIR, and the only difference for different tasks lies in the loss weights and the guidance scale.
The quantitative results are shown in Tab.2. From the Table, one could see that our method achieves comparable even better results than those specific-designed methods. In the zero-shot methods, our method obtains the best and the second performance in terms of SSIM and PSNR metrics, respectively. In PSNR metric, our method outperforms the other methods except ExCNet with the margin of 0.53-5.04. Although ExCNet achieves the better PSNR value, our method achieves the best SSIM values and exceeds it with a margin of 0.046. Besides, even compared to the classic learning-based methods, our method is competitive and obtains better results than some of them. For instance, our method obtains 2.89/0.143 higher PSNR/SSIM value than LNet, 1.43 higher PSNR value than EGAN, and 0.067 higher SSIM value than RNet. Only the MBLLEN could outperform our method on both the PSNR and SSIM metrics. Moreover, we present and analyze the qualitative results in Fig.4, which demonstrate our method enlightens the dark images reliably and reasonably.
@@
We compare our method with four zero-shot image denoising methods, i.e., BM3D [6], N2S [1], N2V [19], and DIP [36]. The above methods except for DIP are specially designed for image denoising. The quantitative results are presented in Tab.3. From the table, one could see that our method obtains comparable even better results than those task-specific methods. To be specific, our method obtains 0.032 and 0.019 higher SSIM value than BM3D and DIP, respectively. Meanwhile, our method also outperforms N2V and DIP in PSNR value with a large margin of 1.00 and 0.59, respectively. Although not achieving the best performance, our method enjoys the appealing capacity of addressing OIR in one model. In addition, we present the qualitative results in Fig.5, from which one could see that our method obtains clearer results with better fidelity to ground truths than the compared methods.

### v4

@@ 3.3 Test-time Degradation Adapter
The essence of OIR lies in addressing the unidentified distribution shifts between the test and training data, while TTA recently emerges as an effective methodology to address this inherent disparities. With the motivations, TDA is presented for adapting PDM to the downstream shifted distribution of degradations during the test phase.
TDA employs a four-layer convolution neural network as the adapter $\phi$, which is optimized for aligning the domain of the generated images $\hat{x}^{g}_{0}$, $\hat{x}^{a}_{0}$ to that of the test degraded image $y$. Note that $\hat{x}^{g}_{0}$ and $\hat{x}^{a}_{0}$ are the estimated outputs from $x^{g}_{t}$ and $x^{a}_{t}$ in terms of Eq.(1), respectively. Specifically, we introduce adversarial training which has shown to be effective in aligning the source and target domains, and the translated samples usually capture rich visual characteristics inherent to the target domain. Therefore, TDA aligns $\phi([\hat{x}^{g}_{0},\hat{x}^{a}_{0}])$ to $y$ which involves significant visual degradations through
$\mathcal{L}_{adv}=\log(1-\mathcal{D}(\phi([\hat{x}^{g}_{0},\hat{x}^{a}_{0}])),$ (5)
@@
$\mathcal{L}_{dis}=-\log\mathcal{D}([y,y])-\log(1-\mathcal{D}(\phi([\hat{x}^{g}_{0},\hat{x}^{a}_{0}])).$ (6)
Meanwhile, TDA adopts a reconstruction loss to maintain the content consistency during the domain alignment, i.e.,
$\mathcal{L}_{rec}=\parallel[\hat{x}^{g}_{0},\hat{x}^{a}_{0}]-\phi([\hat{x}^{g}_{0},\hat{x}^{a}_{0}])\parallel_{2}^{2}.$ (7)
In addition, the perceptual loss is introduced to remain the semantic consistency during the domain alignment, i.e.,
$\mathcal{L}_{pec}=\parallel\mathcal{V}([\hat{x}^{g}_{0},\hat{x}^{a}_{0}])-\mathcal{V}(\phi([\hat{x}^{g}_{0},\hat{x}^{a}_{0}]))\parallel_{2}^{2},$ (8)
@@
where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are the loss weights. By iterative optimization during the denoising process, $\phi$ gradually translates $\hat{x}^{g}_{0}$, $\hat{x}^{a}_{0}$ into the domain of $y$. Note that $\hat{x}^{a}_{0}$ is introduced to prevent $\phi$ from the trivial solution of identity mapping.
@@ 3.4 Adapter-guided Image Restoration
With the adapter $\phi$, $\hat{x}^{g}_{0}$ could be translated into the degraded domain of $y$ through $\phi(\hat{x}^{g}_{0})$, which could be further supervised towards the $y$, and in turn produce the $\mathcal{G}$ to update the generated images $\hat{x}^{g}_{0}$ to the desired clean output $x^{g}_{0}$. AIR is devised to perform this supervised updating process.
Existing studies (Choi et al., 2022) have show that the denoising process exhibits an intriguing temporal dynamic, i.e., $\hat{x}^{g}_{0}$ is generated from the unrecognizable contents to the perceptually rich contents, and finally the imperceptible details. In other words, the contents to be restored are varying as the denoising steps. To take full advantages of PDM, AIR empirically divides the denoising steps into three stages, and guides the restoration through different strategies.
In the first stage where $t\sim T$, the generated images have a lower signal-to-noise ratio, wherein the contents are unrecognizable and only contain some global information such as colors, layouts distributed throughout image. Therefore, AIR applies mse loss on the spatial pixels and color channels to learn global contents, i.e.,
@@
where $\gamma_{1}$ is the loss weight and $\hat{x}_{0}\in\{\hat{x}^{g}_{0},\hat{x}^{a}_{0}\}$. Note that AIR also use this guidance on the auxiliary image $x^{a}_{0}$ so that it involves the similar global properties as $y$, since it is beneficial to the optimization of adapter on most degradations. In addition, a linear warmup is employed on $\gamma_{1}$ to gradually involve this guidance.
In the second stage where $t\sim T/2$, PDM prefers to generate the perceptually rich contents, and thus AIR focuses more on restoring the perceptual contents from $y$. To this end, the perceptual loss is adopted, i.e.,
@@
Meanwhile, an adversarial loss is also introduced to boost the perceptual quality, i.e.,
@@
-- $\log\mathcal{D}([\hat{x}^{g}_{0},\hat{x}^{a}_{0}]-\phi([\hat{x}^{g}_{0},\hat{x}^{a}_{0}])).$
With this loss, $\hat{x}^{g}_{0}$ will be further updated toward the clean image whose degraded version matches $y$, by aligning the domain of content loss caused by degradations. Besides, the mse loss is also employed to remain the global contents, i.e.,
@@
In the third stage where $t\sim 0$, the perceptually rich contents are already presented in $\hat{x}^{g}_{0}$. AIR is only required to recover the imperceptible details. In other words, $\hat{x}^{g}_{0}$ is very close to the natural image, and thus some extra image priors could be introduced to regulate the final output. In consideration of generality, we only introduce a total variation loss, i.e., $\mathcal{G}=\gamma_{5}*TVLoss(\hat{x}^{g}_{0})$, where $\gamma_{5}$ is the loss weight. Note that although the third stage mainly affects the imperceptible details, the pixel values still fluctuate and thus the guidance loss in the second stage is used for a better numerical fidelity.
@@ 4 Experiments
Table 1: Quantitative results of image dehazing on HSTS. Our method outperforms zero-shot methods, and obtains comparable even better results than classic methods through supervised learning. All the compared methods are specifically designed for image dehazing.
@@
PSNR $\uparrow$ 23.96 18.40 19.15 21.69 18.42 15.75 17.41 18.18 18.22 21.24 22.29
SSIM $\uparrow$ 0.902 0.826 0.860 0.868 0.854 0.783 0.833 0.814 0.832 0.835 0.877
Figure 3: Qualitative results on image dehazing, from which one could observe that existing methods excessively dehazing resulting in darkening and/or artifacting of the images. In contrast, our method obtains clearer results which are closer to the natural ground truths.
4.1 Experimental Settings
To evaluate the effectiveness of TAO on addressing OIR, we conduct experiments on multiple tasks of image restoration with exactly the same settings, except for the loss weights and guidance scale ($\lambda_{1-3}$, $\gamma_{1-5}$, $s$), which are finetuned for different types of degradation to obtain the best results. In experiments, we employ an unconditional image diffusion model (Dhariwal & Nichol, 2021) pretrained on ImageNet (Deng et al., 2009) as PDM, and sets the timestep as $T=1000$, which is further divided into three stages in a heuristic way, i.e., the first stage 999-700, the second stage 700-50, and the third stage 50-0. Since PDM is pre-trained on large-scale natural images for high-quality image generation, any degradation in the test image could be regraded as the unforeseen. Both the adapter and discriminators are four-layer convolutional networks, and optimized once at each denoising step through Adam optimizer with default learning rate of 1e-3. First nine layers of pretrained VGG-16 (Simonyan & Zisserman, 2014) network are employed to extract the semantic feature maps for calculating the perceptual loss. For evaluations, considering the specificity of our methods, i.e., adaptation on single test sample, we mainly compare it with those task-specific zero-shot methods, and some classic learning-based methods. To access their performance, PSNR and SSIM metrics are employed, and all experiments are conducted through PyTorch framework on Ubuntu20.04 with GeForce RTX 3090 GPUs.
4.2 Experimental Results
Here, we evaluate our method on the tasks of image dehazing, low-light image enhancement, and image denoising.
Image Dehazing aims to eliminate the haze and boost the visibility of hazy image, which has spawned many methods in the past decades. In experiments, we introduce the HSTS dataset from RESIDE (Li et al., 2018a) for evaluations. To be specific, RESIDE is a large scale haze image dataset, and HSTS is one of the test subsets, which contains 10 synthetic and 10 real-world hazy images. Before the evaluation, we first center crop the images along the shorter edges, and then resize them to match the image size of the PDM.
For a comprehensive comparison, we compare TAO with 10 representative methods which are specifically designed for image dehazing, and could be roughly divided into two groups, i.e., classic learning-based and zero-shot methods. Specifically, classic learning-based methods are DehazeNet (DEN) (Cai et al., 2016), MSCNN (MSN) (Ren et al., 2016), AOD-Net (AOD) (Li et al., 2017), and CAP (Zhu et al., 2015). Zero-shot methods are DCP (He et al., 2010), BCCR (Meng et al., 2013), GRM (Chen et al., 2016), NLD (Berman et al., 2016), DDIP (Gandelsman et al., 2019) and YOLY (Li et al., 2021b).
The quantitative results are presented in Tab.1, from which one could see that our method achieves comparable even better results than those task-specific methods. To be specific, our method obtains the best performance in zero-shot methods, and outperforms YOLY and DCP with 1.05/0.042 and 3.87/0.023 according to PSNR/SSIM metrics, respectively. Meanwhile, even compared with the classic learning-based methods, our method still outperforms most of them and achieves comparable results with the others. For example, our method obtains 3.89/0.051, 3.14/0.017, and 0.60/0.009 higher PSNR/SSIM values than MSN, AOD and CAP, respectively. Although our method is lower than DEN, the performance gaps are not huge, e.g., 1.67/0.025 lower than DEN in terms of PSNR/SSIM metrics. In addition, we also show and analyze the qualitative results in Fig.3, from which one could observe that our method achieves better fidelity and realness than the compared methods.
@@ 4.2 Experimental Results
PSNR $\uparrow$ 13.39 17.87 18.03 14.85 17.15 15.33 15.71 11.24 15.75 17.10
SSIM $\uparrow$ 0.623 0.699 0.809 0.815 0.720 0.763 0.765 0.543 0.665 0.804
Figure 4: Qualitative results on low-light image enhancement, from which one could see that our results are not as smooth as MBLLEN nor as dark as ZDEC. Although there are slight color biases from ground truths, our method achieves a rational lighting of the dark images.
Low-light Image Enhancement aims at improving the perception of the image captured in the environment with poor illumination. To evaluate our method on this task, we introduce the test subset from LOL (Wei et al., 2018b) dataset. LOL includes 485 training and 15 test image pairs of low- and normal-light, which contain noises produced during the photo capture process. Since the images have a resolution of 400$\times$600, we first center crop them along the shorter edges, and then resize them to be applicable to the PDM.
For a comprehensive comparison, we compare our method with nine low-light image enhancement methods, which are divided into two categories, i.e., classical learning-based and zero-shot methods. To be specific, the classical learning-based methods are LightenNet (LNet) (Li et al., 2018b), Retinex-Net (RNet) (Wei et al., 2018a), MBLLEN (Lv et al., 2018), and EnlightenGAN (EGAN) (Jiang et al., 2021). The zero-shot methods are ExCNet (Zhang et al., 2019), Zero-DCE (ZDEC) (Guo et al., 2020), Zero-DCE++ (ZDEC+) (Li et al., 2021c), RRDNet (RRDN) (Zhu et al., 2020), and GDP (Fei et al., 2023). The above methods except GDP are specially designed for low-light image enhancement, while our method is devised for more general purpose of OIR, and the only difference for different tasks lies in the loss weights and the guidance scale.
The quantitative results are shown in Tab.2. From the Table, one could see that our method achieves comparable even better results than those specific-designed methods. In the zero-shot methods, our method achieves the best and the second performance in terms of SSIM and PSNR metrics, respectively. In PSNR, our method outperforms the other methods except ExCNet with the margins of 1.35-5.86. Although ExCNet achieves a slightly better PSNR value of 0.05, our method achieves the best SSIM values and exceeds it with a margin of 0.084. Besides, even compared with the classic learning-based methods, our method is competitive and obtains better results than most of them. For instance, our method obtains 3.71/0.181 higher PSNR/SSIM value than LNet, 2.25 higher PSNR value than EGAN, and 0.105 higher SSIM value than RNet. Only MBLLEN outperforms our method on both the PSNR and SSIM metrics with the slight margins of 0.93 and 0.005, respectively. In addition, we present and analyze the qualitative results in Fig.4, which demonstrate our method enlightens the dark images reliably and reasonably.
@@
PSNR $\uparrow$ 26.35 28.63 28.71 27.17 27.58 28.49
SSIM $\uparrow$ 0.757 0.795 0.864 0.846 0.808 0.830
Figure 6: Visualizations from TDA wherein Auxiliary/Guided Images are translated to Degraded Auxiliary/Guided Images by TDA, whose domains align well to that of Test Degraded Images.
Table 4: Results on different strategies for domain alignment to unknown and unseen degradations. These strategies were implemented based on BasicSR (Wang et al., 2022b). VGAN-S/B denote the Sigmoid Layer plus BCELoss and BCEWithLogitsLoss, respectively.
Metrics Haze Image VGAN_S VGAN_B LSGAN WGAN WGAN_Softplus Hinge
PSNR $\uparrow$ 14.74 22.29 22.06 20.76 13.61 21.47 20.44
SSIM $\uparrow$ 0.770 0.877 0.870 0.873 0.619 0.864 0.860
@@
We compare our method with four zero-shot image denoising methods, i.e., BM3D (Dabov et al., 2007), N2S (Batson & Royer, 2019), N2V (Krull et al., 2019), and DIP (Ulyanov et al., 2018). The above methods except for DIP are specifically designed for image denoising. The quantitative results are presented in Tab.3. From the table, one could see that our method obtains comparable even better results than those task-specific methods. To be specific, our method obtains 0.035 and 0.022 higher SSIM value than BM3D and DIP, respectively. Meanwhile, our method also outperforms N2V and DIP in PSNR value with a large margin of 1.32 and 0.91, respectively. Although not achieving the best performance, our method enjoys the appealing capacity of addressing OIR through one model. In addition, we present the qualitative results in Fig.5, from which one could see that our method obtains clearer results with better fidelity to ground truths than the compared methods.